OpenStudy (anonymous):

find the value of c that makes each trinomial a perfect square. then write the trinomial as a perfect square. the equation , x^2+11x+c

jimthompson5910 (jim_thompson5910):

take half of 11 and square it to get (11/2)^2 = 121/4 This is the value that is missing from x^2+11x to make it a perfect square So c = 121/4 and \[\Large x^2+11x+\frac{121}{4} = \left(x+\frac{11}{2}\right)^2\]

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

so the answer would be 121/4/

jimthompson5910 (jim_thompson5910):

that's the value of c, so that's the first part of the answer

OpenStudy (anonymous):

then what would you do next?

jimthompson5910 (jim_thompson5910):

you mean to factor?

OpenStudy (anonymous):

what would be the trinomial as the perfect square?

jimthompson5910 (jim_thompson5910):

the trinomial as a perfect square is \[\Large \left(x+\frac{11}{2}\right)^2\]

OpenStudy (anonymous):

Ohhh okay i Get it now ! (: Thanks .

jimthompson5910 (jim_thompson5910):

np